A financial analyst wanted to determine the mean annual return on mutual funds. A random sample of 60 returns shows a mean of 12%. If the population standard deviation is assumed to be 4%, estimate with 95% confidence the mean annual return on all mutual funds.


1
Expert's answer
2021-05-14T11:45:22-0400

Given that,


"\bar x = 0.12


,
\sigma = 0.04"",n = 60"




At 95% confidence level the z is ,


"\alpha = 1 - 0.95 = 0.05"


"\dfrac{\alpha }{ 2} = \dfrac{0.05 }{ 2} = 0.025"


"Z_{\frac{\alpha}{2}} = Z_{0.025 }= 1.96"


Margin of error "= E = Z_{\frac{\alpha}{2}}\times \dfrac{\sigma }{\sqrt{n}})"


"=1.96\times \dfrac{0.04}{\sqrt{60}}= 0.010"




At 95% confidence interval estimate of the population mean is,


"\bar x - E < \mu < \bar x + E


\\
0.12 - 0.010 < \mu < 0.12 + 0.010


\\
0.11 < \mu < 0.13


\\
(0.11, 0.13)
\\


(11, 13)"


Confidence interval is (11%,13%)